# Leveraging Modality Tags for Enhanced Cross-Modal Video Retrieval

### v1

@@ III-C Architecture
CLIP4Clip [77] - Neurocomp.’22 44.5 71.4 15.3 197.5 42.1 71.9 16.2 195.4 43.4 70.2 17.5 194.2
CenterCLIP [78] - SIGIR’22 44.2 71.6 15.1 197.9 43.7 71.3 16.9 195.8 - - - -
X-Pool [26] - CVPR’22 46.9 72.8 14.3 201.9 43.9 72.5 14.6 198.7 - - - -
TS2-Net [13] - ECCV’22 47.0 74.5 13.0 205.3 - - - - 41.8 71.6 14.8 195.4
EMCL-Net [15] - NeurIPS’22 46.8 73.1 - 203.0 - - - - 46.8 74.3 12.3 204.2
EMCL-Net [15] - NeurIPS’22 - - - - 45.2 71.4 15.0 197.8 44.1 71.7 15.1 197.8
VoP [27] - CVPR’23 44.6 69.9 16.3 194.8 42.7 68.2 15.9 190.2 46.4 71.9 13.6 199.8
MuMUR [79] - Inf. Retr. J.’23 46.4 72.6 13.9 201.2 44.8 72.0 - 199.3 44.4 74.3 - 201.8
PiDRo [29] - ICCV’23 48.2 74.9 12.6 206.4 - - - - 48.6 75.9 11.8 208.9
HBI [80] - CVPR’23 48.6 74.6 12.0 206.6 - - - - 46.9 74.9 12.1 204.5
DiffusionRet [14] - ICCV’23 49.0 75.2 12.1 206.9 - - - - 46.7 74.7 14.3 204.1
Prompt Switch [81] - ICCV’23 47.8 73.9 14.4 203.9 - - - - - - - -
Cap4Video [48] - CVPR’23 49.3 74.3 12.0 207.4 - - - - 52.0 79.4 10.5 218.9
UCoFiA [82] - ICCV’23 49.4 72.1 12.9 205.0 - - - - 46.5 74.8 13.4 205.7
UCoFiA [82] - ICCV’23 - - - - 45.2 69.6 15.7 194.2 42.1 69.2 16.3 190.4
PAU [83] - NeurIPS’23 48.5 72.7 14.0 203.7 - - - - 48.6 76.0 12.9 209.1
TABLE [10] - AAAI’23 47.1 74.3 13.4 204.3 - - - - 47.9 74.0 14.3 204.0
UATVR [33] - ICCV’23 47.5 73.9 12.3 204.9 - - - - 43.1 71.8 15.1 197.2
TeachCLIP [30] - CVPR’24 46.8 74.3 - - - - - - 43.7 71.2 - -
MV-Adapter [31] - CVPR’24 46.2 73.2 - 202.1 - - - - 44.3 72.1 - 196.9
BiC-Net [84] - TOMM’24 39.4 75.5 - 201.6 32.8 68.2 - 183.4 - - - -
RAP [85] - ACL’24 44.8 71.4 14.4 197.7 - - - - 42.6 70.4 18.0 192.6
@@
QB-Norm [17] QB CVPR’22 47.2 73.0 - 203.2 - - - - 43.3 71.4 - 195.5
DiCoSA [3] QB IJCAI’23 47.5 74.7 13.2 206.0 - - - - 45.7 74.6 11.7 203.8
DiffusionRet [14] QB ICCV’23 48.9 75.2 12.1 207.2 - - - - 48.9 75.5 14.1 207.7
@@
EMCL-Net [15] DSL NeurIPS’22 51.6 78.1 - 215.0 - - - - - - - -
EMCL-Net [15] DSL NeurIPS’22 - - - - - - - - 47.6 73.5 11.9 203.9
TS2-Net [13] DSL ECCV’22 51.1 76.9 11.7 213.6 - - - - 47.4 74.1 12.9 203.9
TABLE [10] DSL AAAI’23 52.3 78.4 11.4 215.9 - - - - 49.1 75.6 14.8 207.6
UATVR [33] DSL ICCV’23 49.8 76.1 12.3 211.4 - - - - - - - -
RAP [85] DSL ACL’24 - - - - - - - - 47.1 74.1 13.9 203.6
@@
RIVRL [86] - TCSVT’22 12.2 26.8 - 74.4 - - - - - - - -
X-Pool [26] - CVPR’22 - - - - 16.1 35.2 67.2 96.2 - - - -
AME-Net [87] - TOMM’22 - - - - - - - - 7.6 21.5 - 61.9
EMCL-Net [15] - NeurIPS’22 13.2 27.3 327.5 75.3 16.4 36.3 69.5 100.0 11.8 28.8 121.6 80.4
RoME [88] - CoRR’22 - - - - - - - - 6.3 16.9 - 48.4
SwAMP [89] - AISTATS’23 - - - - - - - - 9.4 24.9 - 69.6
MuMUR [79] - Inf. Retr. J.’23 - - - - 16.6 37.5 52.7 104.1 - - - -
UCoFiA [82] - ICCV’23 13.0 27.1 323.8 74.8 16.4 37.2 74.6 101.3 12.4 30.4 123.5 83.8
LTME [90] - ICASSP’24 10.6 24.1 - 66.9 - - - - - - - -
BiC-Net [84] - TOMM’24 - - - - - - - - 8.7 23.9 - 66.1
@@
EMCL-Net [15] DSL NeurIPS’22 14.8 29.4 317.3 80.8 18.2 38.2 63.1 105.7 12.6 31.7 104.5 86.8
@@
TABLE II: Comparison with SOTA across all the datasets. --: unreported results. our reproduced results. IS: Inference Strategy. The highest $RSum$ in each block is marked in bold.
IV Experiments
First, we introduce datasets and metrics considered to evaluate MAC-VR in Sec.IV-A. Then, we introduce the implementation details for reproducibility in Sec. IV-B. Sec. IV-C and Sec. IV-D show the comparison with the baseline and SOTA works, and ablation experiments respectively.
IV-A Datasets and Metrics
MSR-VTT [91] is commonly utilised for video retrieval. It comprises of 10K YouTube videos, each with $20$ captions. We utilise two commonly used training splits, with the same test set of 1K videos: the 9k-Train split [5] and the 7k-Train split [70]. DiDeMo [71] collects 10K Flickr videos annotated with paragraph captions. This dataset is evaluated in a video-paragraph retrieval manner [2]. TGIF [72] is a large dataset that contains more than 100K animated short GIFs collected from Tumblr, and natural language sentences annotated via crowdsourcing. We use the split proposed in [92]. Charades [73] contains $9,848$ videos of $157$ action classes in total, where each video is associated with a caption. This dataset consists of fine-grained videos of recorded human actions. We use the standard training and test splits [93]. YouCook2 [74] is a fine-grained dataset that contains 14K YouTube cooking videos that cover $89$ recipes. A textual sentence describes each video clip. Following [70], we evaluate this dataset on the validation set.
Modality-Specific Tags. We extract modality-specific tags from all videos and captions in the datasets above. On average we have between $27$ and $29$ visual and textual tags across all the datasets and splits. Tab. I presents the number of unique visual and textual tags across datasets.
Metrics. We present the retrieval performance for text-to-video retrieval task using standard metrics: Recall at $L=1,5$ ($R@L$) and mean rank ($MeanR$). We report the complete tables including $R@10$ and median rank $MR$ in Sec. B in the appx. We use the sum of cross-modal $R@L$ with $L=1,5,10$ as $RSum$ to rank methods.
IV-B Implementation Details
We build our model using DiCoSA [3] as our baseline. Our architecture employs CLIP’s VIT-B/32 [94] as the image encoder and CLIP’s transformer base as the text encoder to encode captions and visual/textual tags. For tag generation, we use the fine-tuned VideoLLaMA2 [37] as the VLM for visual tags and Llama3.1 [95] as the LLM for textual tags. VideoLLaMA2 runs on $8$ sparsely sampled frames with temperature values $\tau\in{0.7,0.8,0.9,1.0}$ – the same values are used for Llama3.1 – to yield diverse tags. During training, we randomly select $N$ tags, while during inference we use fixed $M$ tags for deterministic results; we set $N=M=12$ for training and evaluating MAC-VR, except when using DSL, where $N=6$ and $M=8$.
Following the implementation details of [3], we optimize the model using an Adam optimizer with linear warm-up, setting initial learning rates of 1e-7 for the text and video encoders and 1e-3 for other modules. Unless specified, we set $\tau_{a}=3$ in Eq. 1, $K=8$, and weight parameters $\alpha_{1}=\alpha_{2}=1$. Our MLP consists of two 256-length linear layers with a ReLU activation in between. We use $N_{v}=12$ frames for all datasets except DiDeMo, where $N_{v}=50$ frames are used for video-paragraph retrieval. The model is trained for $10$ epochs with a batch size of $128$ ($64$ for DiDeMo due to GPU memory) and employs a 4-layer transformer to aggregate frame embeddings.
IV-C Results
First, we present the comparison of MAC-VR against the baseline DiCoSA [3]. Then, we compare MAC-VR with SOTA methods, particularly highlighting the impact of inference strategies on fairness of comparison.
Comparison with Baseline. In Fig. 4, we compare MAC-VR with our Baseline DiCoSA trained using the same training parameters, i.e. batch size, $BS$; same number of frames, $N_{v}$; and when using all inference strategies (IS). In all the datasets we improve $\Delta RSum$ compared to our baseline DiCoSA across all IS – by $+0.9$ for MSR-VTT-9k, $+3.4$ for MSR-VTT-7k, $+2.8$ for DiDeMo, $+9.5$ for Charades, $+3.0$ for TGIF and $+0.6$ for YouCook2 when using the DSL approach as inference strategy. The corresponding tables are shown in Sec. B in the appx.
Comparison with SOTA. In Tab. II, we provide a comprehensive comparison of MAC-VR against all SOTA works. To ensure fairness, we divide methods by their inference strategy and consider three different settings: MAC-VR without any inference strategy, MAC-VR with QB, and MAC-VR with DSL.
$R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
@@ IV-C Results
+TT 52.0 77.6 10.4 215.6
+VT+TT 53.2 77.7 10.0 216.2
TABLE III: Ablation on which modality to extract tags from.
@@
VT TT $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
@@
TABLE IV: Ablation on foundation models. VL: Video-LLaMA. VL2: VideoLLaMA2. L2: Llama2. L3.1: Llama3.1.
@@
TABLE V: Ablation on auxiliary inputs. PG: PEGASUS. L3.1: Llama3.1. VL2: VideoLLaMA2.
(a) Training and Inference.
(b) Inference.
Figure 5: Ablation on varying the number of tags across all inference strategies. $(N-M)$ on the x-axis indicates the number of training tags $N$ and inference tags $M$.
Note that we do not include methods that use extra input modalities during training or that rely on significantly stronger pre-training. For example, in [16, 7, 96, 93], models also train on audio data, so they are not directly comparable to MAC-VR, which only uses video and text. Moreover, VAST [96] includes subtitles (when available) in training as additional modality. This work is also pre-trained on a multi-modal video caption dataset. Similarly, CLIP-ViP [32] uses extra pre-training datasets to improve retrieval performance further. We also do not compare to T-MASS [97] as the official code has been retracted after a discovered data leakage.
Overall, MAC-VR achieves best overall performance when combined with DSL inference across 5 of the 6 datasets – only outperformed by [48] on DiDeMo. Results are consistent using QB inference for MSR-VTT (both splits) and without inference strategies on TGIF, Charades and YouCook2. On MSR-VTT (both splits), results are comparable to SOTA without an inference strategy.
Across 5 of the 6 datasets, MAC-VR shows the best performance for any inference strategy. However, on DiDeMo, its performance is lower than SOTA methods for all inference strategies, even though the DSL approach still gives good results. This difference may be influenced by different training parameters (i.e. $BS$ and $N_{v}$) used in other SOTA [15, 29, 80, 14, 48, 82, 83, 10]. To showcase this, we re-ran two SOTA methods [15, 82] using our training settings, and both performed worse than in their original experiments, i.e. -$\text{-}8.8$ for [15] and -$\text{-}15.3$ for [82], and worse than our MAC-VR, i.e. -$\text{-}0.6$ for [15] and -$\text{-}8.0$ for [82]. Another factor could be that we use the same foundation model settings (only 8 frames) to generate tags for every dataset. In the case of DiDeMo, we extract visual tags from the whole video and textual tags by concatenating individual captions. As a result, many frames may not match any labelled segments, adding irrelevant and detrimental information.
Figure 6: Qualitative results across all the datasets. Left Rank: The ranking results of our baseline without using auxiliary tags. Right Rank: The ranking results of MAC-VR, which incorporates extracted visual (VT) and textual (TT) tags to enhance retrieval. Blue and orange colours highlight meaningful tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.
(c) w/o Modality-Specific Tags.
(d) w/ Modality-Specific Tags.
Figure 7: t-SNE plot of visual and textual concepts on MSR-VTT-9k and Charades with/without using auxiliary modality-specific tags.
IV-D Ablation Studies
We provide ablations on MAC-VR over: the number of tags; architecture design; choice of foundation models; and use of auxiliary captions vs. tags. All ablations are performed on the commonly used MSR-VTT-9k dataset. We provide full tables and additional ablations in the appx. (Secs. B and C respectively).
Num. of Tags in Training and Inference. In Fig. 5, we vary the number of visual/textual tags in training/inference – note we keep the same number of video/text tags. In Fig. 5(a), we adjust only the number of tags in training and use the same value during inference. MAC-VR without IS and with QB increases in performance until reaching the best performance with $12$ tags. However, for DSL, $RSum$ increases until the best performance of $RSum=215.8$ with $6$ tags, then the value begins to drop. When varying the number of tags only in inference, as shown in Fig. 5(b), we observe that $RSum$ increases rapidly and overcomes our baseline DiCoSA when using more than $2$ tags and getting the best performance with $RSum=216.2$ with $8$ tags using DSL. Best performance at inference is reported at $8$ tags with DSL and $12$ tags otherwise, showcasing that using more tags increases performance. The figure also shows that DSL consistently obtains the best performance compared to QB or no inference strategy. Accordingly, we use DSL in all remaining ablation studies.
@@ IV-D Ablation Studies
Choice of Tag Modality In Tab. V, we ablate which modalities to extract tags from for MAC-VR. Using visual and textual tags individually gets results that are similar or slightly superior to our baseline. When combined, all metrics improve, outperforming our baseline.
Choice of Foundation Models. In Tab. V, we ablate the use of different foundation models to extract visual and textual tags from a video and its caption. We compare Video-LLaMA [38] against VideoLLaMA2 [37] and Llama2 [40] against Llama3.1 [95] to extract visual and textual tags respectively. Results show that all metrics improved using VideoLLaMA2 and Llama3.1 by $\Delta RSum=2.2$. Video-LLaMA and Llama2 tend to hallucinate tags more often, qualitative examples can be seen in Sec. D in appx.
Using Auxiliary Captions over Tags. In Tab. V, we show that tags are more informative compared to further captions extracted directly from the video or paraphrased from the caption. We use Blip2 [34]/VideoLLaMA2 [37] as video captioners and PEGASUS [98]/Llama3.1 [95] as paraphrasers, see appx. for details (Sec. E). Results show that using tags outperforms additional captions – including captions from the same models used to generate tags.
V Qualitative Results

### v3

@@ 4.2 Results
CLIP4Clip [Luo et al.(2022a)Luo, Ji, Zhong, Chen, Lei, Duan, and Li] - Neurocomp.’22 44.5 71.4 15.3 197.5 42.1 71.9 16.2 195.4 43.4 70.2 17.5 194.2
CenterCLIP [Zhao et al.(2022)Zhao, Zhu, Wang, and Yang] - SIGIR’22 44.2 71.6 15.1 197.9 43.7 71.3 16.9 195.8 - - - -
X-Pool [Gorti et al.(2022)Gorti, Vouitsis, Ma, Golestan, Volkovs, Garg, and Yu] - CVPR’22 46.9 72.8 14.3 201.9 43.9 72.5 14.6 198.7 - - - -
TS2-Net [Liu et al.(2022)Liu, Xiong, Xu, Cao, and Jin] - ECCV’22 47.0 74.5 13.0 205.3 - - - - 41.8 71.6 14.8 195.4
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] - NeurIPS’22 46.8 73.1 - 203.0 - - - - 46.8 74.3 12.3 204.2
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] - NeurIPS’22 - - - - 45.2 71.4 15.0 197.8 44.1 71.7 15.1 197.8
VoP [Huang et al.(2023)Huang, Gong, Pan, Jiang, Lv, Li, and Wang] - CVPR’23 44.6 69.9 16.3 194.8 42.7 68.2 15.9 190.2 46.4 71.9 13.6 199.8
MuMUR [Madasu et al.(2023)Madasu, Aflalo, Stan, Rosenman, Tseng, Bertasius, and Lal] - Inf. Retr. J.’23 46.4 72.6 13.9 201.2 44.8 72.0 - 199.3 44.4 74.3 - 201.8
PiDRo [Guan et al.(2023)Guan, Pei, Shao, Liu, Li, Gu, Xu, Xu, Yan, and Lam] - ICCV’23 48.2 74.9 12.6 206.4 - - - - 48.6 75.9 11.8 208.9
HBI [Jin et al.(2023a)Jin, Huang, Xiong, Tian, Liu, Ji, Yuan, and Chen] - CVPR’23 48.6 74.6 12.0 206.6 - - - - 46.9 74.9 12.1 204.5
DiffusionRet [Jin et al.(2023c)Jin, Li, Cheng, Li, Ji, Liu, Yuan, and Chen] - ICCV’23 49.0 75.2 12.1 206.9 - - - - 46.7 74.7 14.3 204.1
Prompt Switch [Deng et al.(2023)Deng, Chen, Qin, Chen, and Wu] - ICCV’23 47.8 73.9 14.4 203.9 - - - - - - - -
Cap4Video [Wu et al.(2023a)Wu, Luo, Fang, Wang, and Ouyang] - CVPR’23 49.3 74.3 12.0 207.4 - - - - 52.0 79.4 10.5 218.9
UCoFiA [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal] - ICCV’23 49.4 72.1 12.9 205.0 - - - - 46.5 74.8 13.4 205.7
UCoFiA [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal] - ICCV’23 - - - - 45.2 69.6 15.7 194.2 42.1 69.2 16.3 190.4
PAU [Li et al.(2023a)Li, Song, Gao, Zhu, and Shen] - NeurIPS’23 48.5 72.7 14.0 203.7 - - - - 48.6 76.0 12.9 209.1
TABLE [Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan] - AAAI’23 47.1 74.3 13.4 204.3 - - - - 47.9 74.0 14.3 204.0
UATVR [Fang et al.(2023)Fang, Wu, Liu, Zhou, Song, Wang, Shu, Ji, and Wang] - ICCV’23 47.5 73.9 12.3 204.9 - - - - 43.1 71.8 15.1 197.2
TeachCLIP [Tian et al.(2024)Tian, Zhao, Xin, Lan, and Li] - CVPR’24 46.8 74.3 - - - - - - 43.7 71.2 - -
MV-Adapter [Jin et al.(2024)Jin, Zhang, Gong, Xu, Deng, Wang, Zhang, Shen, and Feng] - CVPR’24 46.2 73.2 - 202.1 - - - - 44.3 72.1 - 196.9
BiC-Net [Han et al.(2024)Han, Zeng, Shi, Xiao, Chen, and Chen] - TOMM’24 39.4 75.5 - 201.6 32.8 68.2 - 183.4 - - - -
RAP [Cao et al.(2024)Cao, Tang, Huang, Jin, Zhang, Liu, Chen, Liang, Yuan, and Li] - ACL’24 44.8 71.4 14.4 197.7 - - - - 42.6 70.4 18.0 192.6
@@
QB-Norm [Bogolin et al.(2022)Bogolin, Croitoru, Jin, Liu, and Albanie] QB CVPR’22 47.2 73.0 - 203.2 - - - - 43.3 71.4 - 195.5
DiCoSA [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] QB IJCAI’23 47.5 74.7 13.2 206.0 - - - - 45.7 74.6 11.7 203.8
DiffusionRet [Jin et al.(2023c)Jin, Li, Cheng, Li, Ji, Liu, Yuan, and Chen] QB ICCV’23 48.9 75.2 12.1 207.2 - - - - 48.9 75.5 14.1 207.7
@@
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] DSL NeurIPS’22 51.6 78.1 - 215.0 - - - - - - - -
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] DSL NeurIPS’22 - - - - - - - - 47.6 73.5 11.9 203.9
TS2-Net [Liu et al.(2022)Liu, Xiong, Xu, Cao, and Jin] DSL ECCV’22 51.1 76.9 11.7 213.6 - - - - 47.4 74.1 12.9 203.9
TABLE [Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan] DSL AAAI’23 52.3 78.4 11.4 215.9 - - - - 49.1 75.6 14.8 207.6
UATVR [Fang et al.(2023)Fang, Wu, Liu, Zhou, Song, Wang, Shu, Ji, and Wang] DSL ICCV’23 49.8 76.1 12.3 211.4 - - - - - - - -
RAP [Cao et al.(2024)Cao, Tang, Huang, Jin, Zhang, Liu, Chen, Liang, Yuan, and Li] DSL ACL’24 - - - - - - - - 47.1 74.1 13.9 203.6
@@
RIVRL [Dong et al.(2022b)Dong, Wang, Chen, Qu, Li, He, and Wang] - TCSVT’22 12.2 26.8 - 74.4 - - - - - - - -
X-Pool [Gorti et al.(2022)Gorti, Vouitsis, Ma, Golestan, Volkovs, Garg, and Yu] - CVPR’22 - - - - 16.1 35.2 67.2 96.2 - - - -
AME-Net [Han et al.(2022)Han, Chen, Zhang, Wang, and Chen] - TOMM’22 - - - - - - - - 7.6 21.5 - 61.9
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] - NeurIPS’22 13.2 27.3 327.5 75.3 16.4 36.3 69.5 100.0 11.8 28.8 121.6 80.4
RoME [Satar et al.(2022)Satar, Zhu, Zhang, and Lim] - CoRR’22 - - - - - - - - 6.3 16.9 - 48.4
SwAMP [Kim(2023)] - AISTATS’23 - - - - - - - - 9.4 24.9 - 69.6
MuMUR [Madasu et al.(2023)Madasu, Aflalo, Stan, Rosenman, Tseng, Bertasius, and Lal] - Inf. Retr. J.’23 - - - - 16.6 37.5 52.7 104.1 - - - -
UCoFiA [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal] - ICCV’23 13.0 27.1 323.8 74.8 16.4 37.2 74.6 101.3 12.4 30.4 123.5 83.8
LTME [Cheng et al.(2024a)Cheng, Kong, Wang, Qu, and Jiang] - ICASSP’24 10.6 24.1 - 66.9 - - - - - - - -
BiC-Net [Han et al.(2024)Han, Zeng, Shi, Xiao, Chen, and Chen] - TOMM’24 - - - - - - - - 8.7 23.9 - 66.1
@@
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] DSL NeurIPS’22 14.8 29.4 317.3 80.8 18.2 38.2 63.1 105.7 12.6 31.7 104.5 86.8
@@
Table 3: Comparison with SOTA across datasets. -: unreported results. our reproduced results. IS: Inference Strategy. The highest performance in each block is marked in bold.
Another factor could be that we use the same foundation model settings (only 8 frames) to generate tags for every dataset. In the case of DiDeMo, we extract visual tags from the whole video and textual tags by concatenating individual captions. As a result, many frames may not match any labelled segments, adding irrelevant and detrimental information. We leave this investigation for future work.
4.3 Ablation Studies
We provide ablations on MAC-VR using MSR-VTT-9k over: the number of tags and architecture design; choice of foundation models; and use of auxiliary captions vs. tags. We provide Full tables and additional ablations in the supp. (Secs. E and F respectively).
Figure 5: Varying number of tags in Training/Inference (left) and Inference (right) across IS. $(n\text{-}m)$ indicates the num. of training tags $n$ and inference tags $m$.
Num. of Tags in Training and Inference. In Fig. 5 (left), we adjust only the number of tags in training and use the same value in inference. MAC-VR without IS and with QB increases in performance until reaching the best one with $12$ tags. However, for DSL, $RSum$ increases until the best performance of $RSum=215.8$ with $6$ tags, then the value drops. When varying the number of tags only in inference, as shown in Fig. 5 (right), $RSum$ increases rapidly and overcomes our baseline DiCoSA when using more than $2$ tags and gets the best performance with $RSum=216.2$ with $8$ tags using DSL.
Method $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
@@ 4.3 Ablation Studies
+$T$ 52.0 77.6 10.4 215.6
+VT+$T$ 53.2 77.7 10.0 216.2
Table 4: Ablation on architecture design.
@@
VT $T$ $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
@@
Table 5: Ablation on foundation models. VL: Video-LLaMA. VL2: VideoLLaMA2. L2: Llama2. L3.1: Llama3.1.
@@
Table 6: Ablation on auxiliary inputs. PG: PEGASUS. L3.1: Llama3.1. VL2: VideoLLaMA2.
Inference Tags $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
(i) original 53.2 77.7 10.0 216.2
(ii) original wrong 50.4 76.5 10.8 211.4
(iii) self-correct 53.9 78.1 9.4 218.1
(iv) self-correct wrong 49.9 76.0 11.1 210.7
Table 7: Ablation on quality of auxiliary tags.
Best performance at inference is reported at $8$ tags with DSL and $12$ tags otherwise, showcasing that using more tags increases performance. The figure also shows that DSL consistently outperforms QB or no IS. Accordingly, we use DSL in other ablation studies.
@@
Choice of Tag Modality. In Tab. 7, we ablate which modalities to extract tags from. Using visual and textual tags individually gets results that are similar or slightly superior to baseline. When combined, all our metrics improve over the baseline.
Choice of Foundation Models. In Tab. 7, we ablate the use of different foundation models to extract visual and textual tags from a video and its caption. We compare Video-LLaMA [Zhang et al.(2023a)Zhang, Li, and Bing] against VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.] and Llama2 [Touvron et al.(2023b)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, et al.] against Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] to extract visual and textual tags respectively. Results show that all metrics improved using VideoLLaMA2 and Llama3.1 by $\Delta RSum=2.2$. Video-LLaMA and Llama2 tend to hallucinate tags more often, qualitative examples can be seen in Sec. G in supp.
Auxiliary Captions vs Tags. Tab. 7 shows that tags are more informative compared to additional captions extracted directly from the video or paraphrased from the caption. We use Blip2 [Li et al.(2023b)Li, Li, Savarese, and Hoi] and VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.] as video captioners and PEGASUS [Zhang et al.(2020)Zhang, Zhao, Saleh, and Liu] and Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] as paraphrasers, see supp. for details (Sec. H). Results show that tags outperforms additional captions – including captions from the same models used to generate tags.
Auxiliary Tags Quality. Tab. 7 evaluates tag quality during inference under four settings: (i) tags extracted from VLMs/LLMs; (ii) incorrect tags obtained by shuffling among samples; (iii) filtered tags, where a VLM/LLM filters the tag set to keep tags relevant to the video/caption (higher quality); and (iv) incorrect, filtered tags obtained by shuffling the filtered tags. Incorrect tags drop $RSum$ by -$\text{-}4.8$ compared to (i), indicating that the auxiliary tags provide semantically relevant information. Using filtered tags improves $RSum$ by $\text{+}1.9$, suggesting that simple post-hoc filtering mitigates VLM/LLM hallucinations at the cost of additional filtering. Finally, wrong-filtered tags perform worst overall ($RSum=210.7$), reinforcing that curated tags are only beneficial when matched to the correct sample. We leave more sophisticated hallucination-mitigation approaches to future work.
5 Qualitative Results